length of each equal side of a triangle is 4m less from the third side.If the perimeter of a triangle is 55m find the dimensions of a triangle.
Answer by mananth(16946)

third side = x
equal side = x-4
..
x+x-4+x-4=55
3x-8=55
3x=55+8
3x=63
x= 21 third side
other sides = 21-4=17
..
17+17+21=55
...
m.ananth@hotmail.ca